# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

### v1

@@ 6 Global modelling: Photometric and RVs
We performed a global fit of all available transit light curves from TESS, MuSCAT2 and MuSCAT3 telescopes, and radial velocity measurements from Subaru-8.2m/IRD spectrograph described in Sections 2, using the Metropolis-Hastings (Metropolis et al. 1953; Hastings 1970) method implemented in TRAFIT, a revised version of Markov-chain Monte Carlo (MCMC) code described in Gillon et al. (2010, 2012, 2014). The RVs collected by Subaru/IRD spectrograph were modeled with a two-body Keplerian model (Murray & Correia 2010), while the transit light curves were modeled using the Mandel & Agol (2002) quadratic limb-darkening model, multiplied by a transit baseline, aiming to correct for external systematic effects (time, airmass, x- and y- position of the star on the detector, background and FWHM of the PSF).
@@
The baseline model for each transit was selected by minimizing the BIC (Bayesian information criterion, Schwarz (1978)), and photometric error bars were re-scaled using the correction factor $CF=\beta_{w}\times\beta_{r}$, where, $\beta_{r}$ is the red noise and $\beta_{w}$ is the white noise (Gillon et al. 2012). The error bars of TOI-6008 RVs measurements were quadratically re-scaled using the ”jitter” noise $=(<Error_{RVs}>^{2}-\sigma^{2}_{RVs})^{1/2}$, where, $<Error_{RVs}>$ is the mean error of the RVs measurements, and $\sigma_{RVs}$ is the standard deviation of the best-fit residuals.
@@
For each system, the jump parameters that were sampled by MCMC: the transit epoch, the planet orbital period, impact parameter, the transit depth, the total transit duration, and the combination $q_{1}=(u_{1}+u_{2})^{2}$ and $q_{2}=0.5u_{1}(u_{1}+u_{2})^{-1}$ (Kipping 2013) of its quadratic limb-darkening coefficients $u_{1}$ $u_{2}$, which are calculated from Claret et al. (2012); Claret (2018).
@@
For each target, we applied a Gaussian prior distribution on the stellar parameters derived from the SED and spectroscopy analysis ($R_{\star}$: stellar radius, $M_{\star}$: stellar mass, $T_{\rm eff}$: stellar effective temperature, $[Fe/H]$: stellar metallicity and $\log g_{\star}$: stellar surface gravity). Two MCMC analyses were performed for each system, the first assuming a circular orbit (i.e $ecc=0$) and the second assuming an eccentric orbit. Based on the Bayes factor $BC=\exp{(-\Delta BIC/2)}$, our results are well compatible with a circular orbit solution.
@@
For each transit light curve, we performed a preliminary analysis composed of 1$\times$×Markov-chain of 100,000 steps to compute the $CF$. The second step, we performed a global MCMC analysis composed of 3$\times$ Markov-chains of 100,000 steps to infer the stellar and planetary physical parameters for each system. We used the Gelman & Rubin (1992) statistical test in order to check for the convergence of each Markov chain.
The results of TOI-5720, TOI-6008 and TOI-6086 are presented in Tables 3, 4 and 5.
@@ 7 Results and discussion
We present the validation and discovery of TOI-5720 b, TOI-6008 b and TOI-6086 b by the TESS mission, which were confirmed using follow-up photometric observations obtained by MuSCAT2, MuSCAT3 and LCOGT-1m0 telescopes (see Figure 3), and radial velocity measurements obtained by Subaru-8.2m/IRD spectrograph (see RVs curve in Figure 6). We characterized the host-stars by combining spectroscopic observations obtained with Shane/Kast (see Section 2.2.4) and IRTF/SpeX (see Section 2.2.5), spectral energy distribution (SED), and stellar evolutionary models (see Section 3.3). We performed a global analysis of TESS data, and ground-based photometric and radial velocity observations in order to constrain the physical parameters for each system (see Section 6). Table 3 shows the stellar properties of host stars TOI-5720, TOI-6008 and TOI-6086. Derived stellar and planetary physical parameters for each system are shown in Tables 3, 5 and 4. Figures 19, 20 and 21 show the parameters posterior distributions for TOI-5720, TOI-6008 and TOI-6086, respectively.
7.1 TOI-5720 b, TOI-6008 b and TOI-6086 b
@@
The host star TOI-5720 is a nearby (23 pc) $K_{\rm mag}=9.54$ M3.5-type star, with an effective temperature of $T_{\rm eff}=3325\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.383\pm 0.018\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.383\pm 0.019\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.19\pm 0.12$ (deduced from the near-infrared spectrum). TOI-5720 b is a short-period Earth-sized planet orbiting around its host star every $P=1.4344555\pm 0.0000036$ days, which has a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming an albedo of 0), an incident flux of $S_{p}=41.7\pm 4.5$ times that of Earth, and an upper limit on the mass of $M_{p}<4\leavevmode\nobreak\ M_{\oplus}$ based on 2 TRES RV measurements with $3\sigma$ error bar (see Section 2.3.2), this confirmed the non-stellar transiting companion. Moreover, the predicted mass using the Chen & Kipping (2017) mass-radius relationship implemented in the forecaster555forecaster: https://github.com/chenjj2/forecaster package, it is found to be $M_{p}=1.32^{+0.92}_{-0.47}$ M.
@@
TOI-6008 is an M5-type ($K_{\rm mag}=9.29$) star at a distance of 36 pc, with an effective temperature of $T_{\rm eff}=3075\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.04$ dex, a radius of $R_{\star}=0.242\pm 0.013\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.230\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.21\pm 0.12$ (deduced from the near-infrared spectrum). TOI-6008 b is a $1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ short-period planet with an orbital period of $P=0.8574347\pm 0.0000424$ day, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K (assuming an albedo of 0) and an incident flux of $S_{p}=41.5\pm 4.5$ times that of Earth. We collected 15 RV measurements of TOI-6008 using the Subaru/IRD spectrograph, the resulting mass is found to be $M_{p}=1.1^{+3.3}_{-0.9}\leavevmode\nobreak\ M_{\oplus}$ with $3\sigma$ error bar. This value ruled out stellar transiting companion, and confirmed a planetary companion nature. Additionally, the predicted planetary mass using the mass-radius relationship implemented in the forecaster package, the value is found to be $M_{p}=1.10^{+0.73}_{-0.36}$ M.
@@
TOI-6086 is a $K_{\rm mag}=9.99$ M3-type star at a distance of 32 pc from the Sun, with an effective temperature of $T_{\rm eff}=3200\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.259\pm 0.013\leavevmode\nobreak\ R_{\odot}$ and a mass of $M_{\star}=0.254\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.2\pm 0.3$ (deduced from the near-infrared spectrum). TOI-6086 b is a short-period ($P=1.3888725\pm 0.0000827$ days) Earth-sized planet ($R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$) orbiting around its host star, with an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K (assuming an albedo of 0), an incident flux of $S_{p}=26.8\pm 2.7$ times that of Earth. Using the mass-radius relationship implemented in the forecaster package, we find that TOI-6086 b has a predicted mass of $M_{p}=1.68^{+1.20}_{-0.55}$ M.
@@
Figures 12 and 13 show the distribution of our planets versus orbital period and equilibrium temperature. These systems are suitable targets for high precision radial velocity facilities in order to provide the planetary mass measurements, bulk densities as well as to search for additional outer planets in the system. Assuming a rocky composition for these planets, the predicted RV semi-amplitudes using the mass-radius relationship by Chen & Kipping (2017), the values are found to be $K_{\rm TOI-5720}=1.4^{+1.0}_{-0.5}$ m/s for TOI-5720, $K_{\rm TOI-6008}=2.0^{+1.3}_{-0.6}$ m/s for TOI-6008 and $K_{\rm TOI-6086}=2.4^{+1.6}_{-0.8}$ m/s for TOI-6086. By combining the predicted RV semi-amplitudes with the relative faintness of the host stars ($V_{\rm mag}$ of 13.9 to 16.0), would require a high precision spectroscopic observations using stabilized spectrographs such as MAROON-X (Seifahrt et al. 2018) mounted on the 8.1m-Gemini-North telescope. Assuming an exposure times of 1800s and good observing conditions, the MAROON-X exposure time calculator (ETC) predict RV precision of 0.6 m/s, 1.1 m/s and 1.0 m/s for TOI-5720, TOI-6008 and TOI-6086, respectively. Using a few RV measurements for each target will allow us to measure the planetary masses, bulk densities and the eccentricities of the orbits with 10-$\sigma$ precision.
@@
The combination of the equilibrium temperature, the large stellar incident flux, the infrared brightness and small sizes of the host stars, making all three planets suitable for transmission and emission spectroscopy studies with JWST. We calculated the transmission spectroscopy metrics (TSM) using equation #1 from Kempton et al. (2018). Our predicted TSM are found to be $9.7^{+7.1}_{-4.1}$ for TOI-5720 b, $18.8^{+12.6}_{-6.8}$ for TOI-6008 b and $12.0^{+8.5}_{-4.7}$ for TOI-6086 b. We also calculated their emission spectroscopy metrics (ESM) using equation #4 from Kempton et al. (2018), and we find that ESM to be $3.4\pm 0.5$ for TOI-5720 b, $6.4\pm 0.7$ for TOI-6008 b and $4.2\pm 0.7$ for TOI-6086 b. Figure 14 shows the planets positions in the TSM and ESM vs planetary radius diagrams of Earth-sized ($R_{p}<1.5$ R) transiting exoplanets around M-dwarfs ($T_{\rm eff}<3800$) observed or scheduled on JWST. Based of TSM and ESM parameters, TOI-6008 b stands out as one of the favourable terrestrial planet for atmospheric characterization (emission, transmission and phase curve photometry) compared to other systems that have been observed with JWST (e.g. TRAPPIST-1 b, TRAPPIST-1 c, TRAPPIST-1 e and TRAPPIST-1 g; Gillon et al. (2017), L 98-59 b; Kostov et al. (2019), GJ 357 b; Luque et al. (2019), GJ 3473 b; Kemmer et al. (2020), LHS 1140 c; Dittmann et al. (2017), LTT 1445A b; Winters et al. (2019), LTT 1445A c; Winters et al. (2022), TOI-1468 b; Chaturvedi et al. (2022), LHS 1478 b; Soto et al. (2021), TOI-732 b; Nowak et al. (2020), LP 791-18 b; Crossfield et al. (2019)). Thanks to the relative high equilibrium temperatures of the planets, small sizes and the brightness of the stars, making also TOI-5720 b and TOI-6086 b good targets for the emission spectroscopy and phase curve photometry compared with TRAPPIST-1 b and TRAPPIST-1 c.
@@
Figure 12: Planetary radius as a function of the orbital period diagram of known transiting exoplanets from NASA Archive of Exoplanets. Grey circles show the FGK planetary systems, and colored points show M-dwarfs systems. M-dwarfs planets are colored by the host star effective temperature. The size of the points scale according to the stellar radius. All short-period ($<$1.5 days) Earth-sized planets orbiting around M-dwarfs are highlighted with error bars, and planets with and without mass measurements are highlighted with + and -, respectively. TOI-5720 b, TOI-6008 b and TOI-6086 b are circled in blue, purple and green, respectively.
Figure 13: Planetary radius as a function of planetary equilibrium temperature diagram of known transiting exoplanets from NASA Archive of Exoplanets. Grey circles show the FGK planetary systems, and colored points show M-dwarfs systems. M-dwarfs planets are colored by the host star effective temperature. The size of the points scale according to the stellar radius. TOI-5720 b, TOI-6008 b and TOI-6086 b are circled in blue, purple and green, respectively.

### v2

@@ 6 Global modeling: Photometry and radial velocities
We performed a global fit of all available transit light curves from the TESS, MuSCAT2, and MuSCAT3 telescopes and of all radial velocity measurements from the Subaru-8.2m/IRD spectrograph described in Sections 2, using the Metropolis-Hastings (Metropolis et al. 1953; Hastings 1970) method implemented in TRAFIT, a revised version of the Markov chain Monte Carlo (MCMC) code described in Gillon et al. (2010, 2012, 2014). The RVs collected by the Subaru/IRD spectrograph were modeled with a two-body Keplerian model (Murray & Correia 2010), while the transit light curves were modeled using the Mandel & Agol (2002) quadratic limb-darkening model, multiplied by a transit baseline, with the aim to correct for external systematic effects (time, airmass, x- and y- position of the star on the detector, background, and FWHM of the PSF).
@@
The baseline model for each transit was selected by minimizing the Bayesian information criterion (BIC; Schwarz (1978)), and photometric error bars were rescaled using the correction factor $CF=\beta_{w}\times\beta_{r}$, where $\beta_{r}$ is the red noise and $\beta_{w}$ is the white noise (Gillon et al. 2012). The error bars of TOI-6008 RV measurements were quadratically rescaled using the jitter noise $=(<Error_{RVs}>^{2}-\sigma^{2}_{RVs})^{1/2}$, where $<Error_{RVs}>$ is the mean error of the RV measurements, and $\sigma_{RVs}$ is the standard deviation of the best-fit residuals.
@@
For each system, the jump parameters were sampled by the MCMC: the transit epoch, the planet orbital period, the impact parameter, the transit depth, the total transit duration, and the combination $q_{1}=(u_{1}+u_{2})^{2}$ and $q_{2}=0.5u_{1}(u_{1}+u_{2})^{-1}$ (Kipping 2013) of its quadratic limb-darkening coefficients $u_{1}$ and $u_{2}$, were calculated from Claret et al. (2012); Claret (2018).
@@
For each target, we applied a Gaussian prior distribution on the stellar parameters derived from the SED and spectroscopy analysis ($R_{\star}$: stellar radius, $M_{\star}$: stellar mass, $T_{\rm eff}$: stellar effective temperature, $[Fe/H]$: stellar metallicity, and $\log g_{\star}$: stellar surface gravity). Two MCMC analyses were performed for each system, the first assuming a circular orbit (i.e., $ecc=0$), and the second assuming an eccentric orbit. Based on the Bayes factor $BC=\exp{(-\Delta BIC/2)}$, our results are well compatible with a circular orbit solution.
@@
For each transit light curve, we performed a preliminary analysis composed of one Markov chain with 100,000 steps to compute the $CF$. In the second step, we performed a global MCMC analysis composed of three Markov chains with 100,000 steps to infer the stellar and planetary physical parameters for each system. We used the Gelman & Rubin (1992) statistical test in order to confirm the convergence of each Markov chain.
The results for TOI-5720, TOI-6008, and TOI-6086 are presented in Tables 3, 4, and 5.
@@ 7 Results and discussion
We present the validation and discovery of TOI-5720 b, TOI-6008 b and TOI-6086 b, by the TESS mission, which were confirmed using follow-up photometric observations obtained with the MuSCAT2, MuSCAT3, and LCOGT-1m0 telescopes (see Figure 3), and radial velocity measurements obtained with the Subaru-8.2m/IRD spectrograph (see RVs curve in Figure 6). We characterized the host-stars by combining spectroscopic observations obtained with Shane/Kast (see Section 2.2.4) and IRTF/SpeX (see Section 2.2.5), the spectral energy distribution (SED), and stellar evolutionary models (see Section 3.3). We performed a global analysis of the TESS data and ground-based photometric and radial velocity observations in order to constrain the physical parameters for each system (see Section 6). Table 3 shows the stellar properties of the host stars TOI-5720, TOI-6008, and TOI-6086. The derived stellar and planetary physical parameters for each system are shown in Tables 3, 5, and 4. Figures 18, 19, and 20 show the posterior distribution parameters for TOI-5720, TOI-6008, and TOI-6086, respectively.
7.1 TOI-5720 b, TOI-6008 b, and TOI-6086 b
@@
The host star TOI-5720 is a nearby (36 pc) $K_{\rm mag}=9.54$ M3.5-type star, with an effective temperature of $T_{\rm eff}=3325\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.383\pm 0.018\leavevmode\nobreak\ R_{\odot}$, a mass of $M_{\star}=0.383\pm 0.019\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.19\pm 0.12$ (deduced from the near-infrared spectrum). TOI-5720 b is a short-period Earth-sized planet orbiting its host star every $P=1.4344555\pm 0.0000036$ days, which has a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming an albedo of 0), an incident flux of $S_{p}=41.7\pm 4.5$ times that of Earth, and an upper limit on the mass of $M_{p}<4\leavevmode\nobreak\ M_{\oplus}$ based on two TRES RV measurements with a $3\sigma$ error bar (see Section 2.3.2). This confirmed the nonstellar transiting companion. Moreover, the predicted mass using the Chen & Kipping (2017) mass-radius relation implemented in the forecaster555forecaster: https://github.com/chenjj2/forecaster package it is found to be $M_{p}=1.32^{+0.92}_{-0.47}$ M.
@@
TOI-6008 is an M5-type ($K_{\rm mag}=9.29$) star at a distance of 23 pc, with an effective temperature of $T_{\rm eff}=3075\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.04$ dex, a radius of $R_{\star}=0.242\pm 0.013\leavevmode\nobreak\ R_{\odot}$, a mass of $M_{\star}=0.230\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.21\pm 0.12$ (deduced from the near-infrared spectrum). TOI-6008 b is a $1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ short-period planet with an orbital period of $P=0.8574347\pm 0.0000424$ day, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K (assuming an albedo of 0), and an incident flux of $S_{p}=41.5\pm 4.5$ times that of Earth. We collected 15 RV measurements of TOI-6008 using the Subaru/IRD spectrograph. The resulting mass is found to be $M_{p}=1.1^{+3.3}_{-0.9}\leavevmode\nobreak\ M_{\oplus}$ with a $3\sigma$ error bar. This value ruled out a stellar transiting companion, and confirmed the nature of the planetary companion. Additionally, the predicted planetary mass using the mass-radius relation implemented in the forecaster package, is found to be $M_{p}=1.10^{+0.73}_{-0.36}$ M.
@@
TOI-6086 is a $K_{\rm mag}=9.99$ M3-type star at a distance of 32 pc from the Sun, with an effective temperature of $T_{\rm eff}=3200\pm 75$ K, a surface gravity of $\log g_{\star}=5.01\pm 0.02$ dex, a radius of $R_{\star}=0.259\pm 0.013\leavevmode\nobreak\ R_{\odot}$, a mass of $M_{\star}=0.254\pm 0.011\leavevmode\nobreak\ M_{\odot}$ (derived from the SED analysis), and a metallicity of $[Fe/H]=-0.2\pm 0.3$ (deduced from the near-infrared spectrum). TOI-6086 b is a short-period ($P=1.3888725\pm 0.0000827$ days) Earth-sized planet ($R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$) orbiting its host star, with an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K (assuming an albedo of 0), an incident flux of $S_{p}=26.8\pm 2.7$ times that of Earth. Using the mass-radius relation implemented in the forecaster package, we find that TOI-6086 b has a predicted mass of $M_{p}=1.68^{+1.20}_{-0.55}$ M.
@@
Figures 12 shows the distribution of our planets versus orbital period and equilibrium temperature. These systems are suitable targets for high-precision radial velocity facilities in order to provide the planetary mass measurements, and bulk densities and to search for additional outer planets in the system. When we assume a rocky composition for these planets, the predicted RV semi-amplitudes using the mass-radius relation by Chen & Kipping (2017), are found to be $K_{\rm TOI-5720}=1.4^{+1.0}_{-0.5}$ m/s for TOI-5720, $K_{\rm TOI-6008}=2.0^{+1.3}_{-0.6}$ m/s for TOI-6008, and $K_{\rm TOI-6086}=2.4^{+1.6}_{-0.8}$ m/s for TOI-6086. A combination of the predicted RV semi-amplitudes with the relative faintness of the host stars ($V_{\rm mag}$ of 13.9 to 16.0) would require high-precision spectroscopic observations using stabilized spectrographs such as MAROON-X (Seifahrt et al. 2018) mounted on the 8.1m Gemini-North telescope. For exposure times of 1800s and good observing conditions, the MAROON-X exposure-time calculator (ETC) predicts an RV precision of 0.6 m/s, 1.1 m/s, and 1.0 m/s for TOI-5720, TOI-6008, and TOI-6086, respectively. A few RV measurements for each target will allow us to measure the planetary masses, bulk densities, and the eccentricities of the orbits with a precision 10 $\sigma$.
@@
The combination of the equilibrium temperature, the high incident flux, the infrared brightness, and small sizes of the host stars, make all three planets suitable for transmission and emission spectroscopy studies with JWST. We calculated the transmission spectroscopy metrics (TSM) using equation 1 from Kempton et al. (2018). Our predicted TSM are $9.7^{+7.1}_{-4.1}$ for TOI-5720 b, $18.8^{+12.6}_{-6.8}$ for TOI-6008 b, and $12.0^{+8.5}_{-4.7}$ for TOI-6086 b. We also calculated their emission spectroscopy metrics (ESM) using equation 4 from Kempton et al. (2018), and we find that ESM to be $3.4\pm 0.5$ for TOI-5720 b, $6.4\pm 0.7$ for TOI-6008 b, and $4.2\pm 0.7$ for TOI-6086 b. Figure 13 shows the planet positions of TSM and ESM versus planetary radius diagrams of Earth-sized ($R_{p}<1.5$ R) transiting exoplanets around M dwarfs ($T_{\rm eff}<3800$) observed or scheduled on JWST. Based on the TSM and ESM parameters, TOI-6008 b is one of favourable terrestrial planet for atmospheric characterization (emission, transmission, and phase curve photometry) compared to other systems that have been observed with JWST (e.g., TRAPPIST-1 b, TRAPPIST-1 c, TRAPPIST-1 e, and TRAPPIST-1 g; Gillon et al. (2017), L 98-59 b; Kostov et al. (2019), GJ 357 b; Luque et al. (2019), GJ 3473 b; Kemmer et al. (2020), LHS 1140 c; Dittmann et al. (2017), LTT 1445A b; Winters et al. (2019), LTT 1445A c; Winters et al. (2022), TOI-1468 b; Chaturvedi et al. (2022), LHS 1478 b; Soto et al. (2021), TOI-732 b; Nowak et al. (2020), LP 791-18 b; Crossfield et al. (2019)). The relatively high equilibrium temperatures of the planets, their small sizes, and the brightness of the stars, also make TOI-5720 b and TOI-6086 b good targets for emission spectroscopy and phase-curve photometry compared with TRAPPIST-1 b and TRAPPIST-1 c.
@@
Figure 12: Planetary radius vs. orbital period (top panel) and Planetary radius vs. planetary equilibrium temperature (bottom panel) diagrams of known transiting exoplanets from NASA Archive of Exoplanets. Grey circles show the FGK planetary systems, and colored points show M-dwarfs systems. M-dwarfs planets are colored by the host star effective temperature. The size of the points scale according to the stellar radius. All short-period ($<$1.5 days) Earth-sized planets orbiting around M-dwarfs are highlighted with error bars, and planets with and without mass measurements are highlighted with + and -, respectively. TOI-5720 b, TOI-6008 b and TOI-6086 b are circled in blue, purple and green, respectively.